# Evolving from Single-modal to Multi-modal Facial Deepfake Detection: A Survey

### v1

@@ 3.1.2 Proactive Detection Methods
Proactive detection methods for deepfakes utilize strategies that preemptively counteract deepfake manipulation before it occurs or facilitate the identification of manipulated content afterward. Unlike passive methods, which assess content post-creation, proactive approaches embed specific signals, such as adversarial perturbations, into the original content. These signals aid in verifying authenticity and integrity if the content is later altered. For effectiveness, these alterations visually resemble the original to evade detection while impairing the functionality of deepfake generation models. In the following subsections, we explore recent proactive detection efforts targeting deepfakes generated by both GANs and DMs.
@@ 3.1.2.1   Proactive Methods for GANs
To further improve the resilience of perturbations, Huang et al. [137] introduce a poison perturbation generator that improves the effectiveness of perturbations using an alternating training strategy. Wang et al. [138] focus on improving the resilience of perturbations by targeting the Lab color space, particularly the decorrelated a and b channels. Expanding defensive strategies, Anejaet al. [139] develop a data-driven method that quickly produces tailored perturbations to redirect manipulation attempts toward a predefined outcome. Huang et al. [140] design the Cross-Model Universal Adversarial Watermark (CMUA-Watermark), broadening defense across various facial images and deepfake models. Asnani et al. [141] and [109] propose using real image templates to improve the detection accuracy of manipulation, helping differentiate between real and manipulated images.
Recent innovations have suggested conducting source tracing alongside injecting adversarial perturbations into the original input to enhance traceability against manipulations. Innovations by Wang et al. [142] and Wu et al. [143] underscore the importance of tracing original source images to boost detection system credibility. Yu et al. [14] enhance traceability by embedding traceable fingerprints directly into the generator’s parameters, linking outputs uniquely to their creators. Addressing gaps in proactive defenses, Tang et al. [144] introduce the Feature-Output ensemble UNiversal Disruptor (FOUND), which disrupts in the latent space and employs a gradient-ensemble algorithm to provide a universal attack pattern against multiple deepfake models, using minimal original-attribute images. In [145], Zhang et al. introduce Dual Defense, which embeds an invisible watermark in target faces to disrupt malicious face swapping while preserving traceability. Their approach utilizes a watermark embedding network that balances invisibility, adversariality, and traceability through perceptual adversarial encoding strategies. In a recent study, Wu et al. [146] demonstrate that traditional watermarking methods, when applied to AI-generated Deepfake content, can interfere with forgery detection. To address this issue, they propose ’AdvMark’, a plug-and-play solution that transforms robust watermarking systems into adversarial frameworks, retaining the ability to extract watermarks for provenance tracking.
@@ 3.1.2.2   Proactive Methods for DMs
Expanding the scope, Zhao et al. [149] examine the effectiveness of watermarking within diffusion models, with Cao et al. [150] extending proactive security measures to audio models. To enhance model transferability, Shim et al. [151] disrupt the latent encoding phase of deepfake generation using LEAT, which employs a Normalized Gradient Ensemble strategy to optimize disruptions across various models.
@@ 3.2 Multi-modal Deepfake Detection
The concurrence of fake content in multiple modalities leads to more convincing deepfakes compared to those visually manipulated ones. In such context, it becomes imperative to develop the capability to detect these intricate multi-modal deepfakes. In this subsection, we provide a comprehensive investigation of the more intricate and challenging multi-modal scenario in this subsection, where auditory or textual content can also be falsified together with the presence of visual manipulation.
3.2.1 Audio-Visual
Audio-visual deepfakes involve the manipulation of either or both the audio and visual streams in a video. The combination of manipulations on these two modalities potentially raises more serious ethical and societal concerns, e.g., manipulating politicians to make false claims convincingly by delivering a lip-synced video with fake audio.
The heterogeneous nature and the modality gap between audio and video make multi-modal audio-visual deepfake detection a more challenging task. Existing explorations on this task commonly design two-stream framework to incorporate multi-modal inputs. These efforts can be categorized into three main paradigms: independent learning, joint learning and matching-based learning.
@@ 3.2.1 Audio-Visual
As a baseline method presented in [10], two independent streams of CNN are proposed for audio and visual deepfake detection separately, and the final multi-modal detection result is obtained by combining the audio and visual predictions. In a preliminary attempt to evaluate the FakeAVCeleb[5] benchmark, Khalid et al.[11] ensemble predictions from independently learned uni-modal detectors. Furthermore, Hashmi et al.[152] devise an audio-visual network besides uni-modal networks and adopt ensemble learning to make the final decision based on predictions from audio, video, and audio-visual networks. Similarly, AVFakeNet[153] leverages Swin-Transformer [175] as backbones for both audio and visual branches for feature extraction and then integrates the result from both branches for the multi-modal prediction.
@@
Despite its intuitive nature, the paradigm of independent learning might overlook the intrinsic relationships between audio and visual data, which can be critical in accurately detecting deepfakes.
@@
Joint Learning This approach integrates audio and visual modalities to leverage multi-modal information for enhanced deepfake detection. The general network architecture of the joint learning paradigm encompasses two modal-specific branches taking as input corresponding modality content for feature extraction, i.e., video frames and audio waveform, and a fusion module to integrate features of different modalities for audio-visual detection. This paradigm mines the inter-modal relationship by feature fusion to enhance the multi-modal deepfake detection. Based on at which stage the modal-specific feature are fused together, joint learning-based detection algorithms can be further divided into intermediate and late fusion manners. Moreover, some existing late fusion-based audio-visual detectors leverage multi-task strategy or impose regularization to facilitate the learned multi-modal representations.
@@
Intermediate Fusion: This method holds the assumption that simply combining the high-level abstract features extracted from both streams are insufficient to mine the inter-modal information. Zhou et al.[10] initially propose the joint audio-visual deepfake detection, handling cases in which either or both visual and auditory modalities have been manipulated. In their method, a sync-stream is designed to leverage the synchronization patterns in addition to separate branches for each modality, which fuses the intermediate features from distinct modalities by cross-attention mechanism. Yang et al.[42] employ a Transformer architecture, where audio and visual inputs go through a temporal-spatial encoder, and corresponding features of both modalities interact with each other via a decoder with bidirectional cross-attention module. Wang et al.[154] incorporate inspiration from how the biological system perceive the multi-modal environment, proposing visual and audio branches as different sensory systems, of which the obtained signals are then integrated by audio-visual attention mechanism acting like human multi-sensory neurons. Specifically, the intermediate audio and visual features attend to each other at all stages and are concatenated at the last few stages for further multi-modal detection.
@@
Late Fusion: In this strategy, the audio and visual representations are only merged in the final stages of feature extraction. Techniques such as concatenation or addition are commonly used to merge modalities in late fusion [11][152][155][156][157][35][13]. Various advanced strategies, such as attention mechanism[158][159] and MLP mixer layer[12] are applied to enhance effective fusion.
@@
Further, Liu et al. [157] propose a tri-branch framework to handle audio, video, and multi-scale frame inputs, where the extracted features are weighted based on inter-modal interactions and concatenated. Moreover, this method also designs Forgery Clues Magnification Transformer for feature extraction of each branch, where the tokens in the input sequence interact with each other to obtain the weighted sequence to magnify the intra-modal inconsistency, and Distribution Difference based Inconsistency Computing module to align information of audio and video modalities.
Dynamic weighted addition has proven effective in fusion strategies for multi-modal tasks. While addressing the temporal audio-visual deepfake localization task, Cai et al.[35] calculate the weighted sum of uni-modal boundary maps from corresponding branches as the fused boundary map, where the weights are dependent on uni-modal boundary maps themselves. Similarly, Wang et al.[13] leverage dynamic weighted averaging as late fusion approach in the proposed AVT2-DWF framework. Using a two-stream transformer architecture for uni-modal feature extraction, the fused information for multi-modal detection is obtained by weighted averaging the classification tokens in both uni-modal features. Notably, the weights AVT2-DWF are calculated by inter-modal attention mechanism.
More sophisticated than simple concatenation or addition, attention mechanisms and MLP mixer layers offer advanced alternatives for multi-modal feature fusion. Sree et al.[159] employ a multi-head self-attention module after concatenating features of different modalities. The attention mechanism further enhances the inter-modal interaction and benefits subsequent multi-modal detection. More than attention mechanism, Zou et al.[158] feed the concatenated audio-visual features into Transformer layers to fuse information from both modalities. Alternatively, Raza et al.[12] employ an MLP mixer layer that not only enhances intra-modal features but also facilitates the integration of inter-modal information by passing concatenated modal-specific features through an inter-modality MLP mixer layer.
@@
The networks proposed in [10, 42] are tasked with three binary classification objectives: uni-modal audio and visual deepfake detection based on learned representations from corresponding branches and multi-modal detection based on the fused signals. Besides simultaneously introducing uni-modal detection tasks, Muppalla et al.[156] make the multi-modal detection more fine-grained, changing the original binary classification to a four-class one. Similarly, Raza et al.[12] reformulate multi-modal audio-visual deepfake detection as a multi-label classification task. They modify the classification head to incorporate fused signals and simultaneously perform classification tasks for both modalities. The multi-task strategy employed in these studies requires more fine-grained labels of each modality to provide supervision during training.
@@
Unlike those incorporating fine-grained uni-modal detection tasks, Yu et al.[160] design an auxiliary task from a disparate perspective. Critiquing the reliance solely on audio-visual synchronization as insufficient for generalizable audio-visual deepfake detection, the paper emphasizes learning content-based cross-modal correlations. In this work, Yu et al. introduce a correlation distillation task, where well-trained audio and visual speech recognition models serve as teachers to obtain content-based cross-modal correlation, and the student detection network learns to model cross-modal correlation from teachers. This explicitly learned cross-modal correlation based on content helps to prevent overfitting.
@@
Enhanced Representation: This line of work aims to learn more discriminative representations for effective detection during training. To this end, various strategies including regularization[42][157][158] and disentanglement[159] are employed.
@@
Yang et al.[42] apply a contrastive loss and an additive angular margin loss to make the fused multi-modal features of manipulated and genuine samples more discriminative. Likewise, Zou et al.[158] also impose the contrastive constraint in cross-modal cases, thus the audio and visual features become more similar when they share common authenticity, i.e., both streams are real. Besides, Zou et al.[158] further introduce within-modal regularization to align uni-modal features with their respective targets, ensuring the uni-modal features are discriminative for corresponding uni-modal detection. The Distribution Difference based Inconsistency Computing module in [157] also leverages contrastive loss to impose regularization on the relationship between audio and video features based on Jensen-Shannon divergence.
@@
In addition, Mis-AvoiDD[159] enhances representations for effective multi-modal deepfake detection by decoupling them into modal-invariant and modal-specific subspaces. Specifically, a modal-invariant encoder and two modal-specific encoders are devised to extract common and specific features from each modality. In modal-invariant subspace, the discrepancy between representations of different modalities is minimized to align modal-specific features. Furthermore, an orthogonal loss imposes constraints on the relationship between the modal-invariant and modal-specific features of the same modality as well as the relationship between the modal-specific features from different modalities. The orthogonal regularization ensures the distinctness between representation subspaces.
@@
Matching-based Learning Distinct from fusion-based learning, matching-based learning leverages the natural synchronization found in genuine audio-visual content, which is typically absent in synthetic pairs[162, 161]. This approach focuses on matching audio and visual signals by evaluating the consistency between extracted features from each modality, rather than using traditional classification techniques. In this paradigm, networks employ modality-specific feature extractors and are trained on authentic content via contrastive learning, enhancing their generalization capabilities across various types of manipulation.
Cheng et al.[162] propose Voice-Face matching detection based on the observation verified in human perception and neurology field, which reveals the homogeneity between the face and voice of one person. The detection network involves only voice and face feature extractors, without fusion module and classification head. The predicted genuineness of the audio-visual pair is determined by the consistency of features from these two modalities. While training the feature extractors, contrastive InfoNCE[176] loss is applied to increase the similarity between matched voice and face feature pairs.
@@
Tian et al.[161] adopt a matching-based Intra-Cross-Consistency method for multi-modal detection, where a matching score from the detection model is compared with a threshold to obtain detection results. Similar to [162], this approach also evaluates cross-consistency between audio and visual features to compute the matching score. The method addresses the trade-off observed in deepfake generation between temporal identity consistency and the synchronization of face motion with audio-visual elements by incorporating temporal identity intra-consistency. This method merely required self-supervised learning on real videos by a contrastive learning strategy, thus becoming generalizable.
@@
Feng et al[165] handle the task from a novel perspective of anomaly detection, requiring training merely on real videos in a self-supervised manner. At the first stage of this method, an audio-visual network is trained to estimate the inter-modal synchronization score, which is used to extract heuristically designed features. Subsequently, an auto-regressive model is then proposed to learn the distribution of previously extracted feature sets. While performing detection, the samples with low likelihood estimated by the auto-regressive model are flagged as fake.
POI-Forensics[163] focuses on identity consistency instead of cross-modal inconsistency. Cozzolinno et al.[163] argue that any manipulation inevitably perturbs characteristics of the portrayed identity. Therefore, the authors utilize the indirect clue that the individual in the video is not consistent with the person who is claimed to be. While performing detection, this method requires a reference set of at least 10 videos of the person of interest, and the similarity index between the target video and reference videos would be used to determine prediction.
Haq et al.[164] introduce novel biological signal emotion psychology to improve deepfake detection. Based on the emotion extracted from each modality, the proposed neurosymbolic framework performs intra-modal emotion transition analysis and inter-modal reasoning. The former aims to explore the temporal emotion inconsistency, and the latter explores the cross-modal emotion consistency according to psychological theory. Notably, the proposed framework does not require training on deepfake datasets owing to the underlying psychology and emotion knowledge, leading to better generalizability.
Yu et al.[166] detect audio-visual deepfakes based on two augmented views of visual signals, proposing a two-stage framework PVASS-MDD to better capture the audio-visual inconsistency. The framework encompasses an auxiliary Predictive Visual-audio Alignment Self-supervision (PVASS) stage as well as the Multi-modal Deepfake Detection (MDD) stage. In the PVASS stage, the visual signals are augmented to two different views, each of which interacts with audio information via the Predictive VA Alignment module to obtain further corresponding VA representations, the self-supervised learning objective is to make the two VA representations similar. In the MDD stage, the two VA representations corresponding to two augmented visual views are concatenated to be fed into a classifier for multi-modal prediction. Besides cross-entropy loss, regularization terms including a contrastive loss and an auxiliary loss are additionally applied.
3.2.2 Text-Visual
Recent advances in text-visual deepfake detection have introduced sophisticated methods to address manipulations on text and visual modalities. Figure 5 illustrates the framework for this multi-modal detection and grounding task.
Figure 5: HAMMER[6] for text-visual deepfake detection and grounding. Figure from [6]

### v3

@@ 3.1.2 Proactive Detection Methods
Proactive detection methods for deepfakes utilize strategies that preemptively counteract deepfake manipulation before it occurs or facilitate the identification of manipulated content afterward. Figure 4 showcases how proactive detection approaches distinguish themselves from passive detection methods. Unlike passive methods, which assess content post-creation, proactive approaches embed specific signals, such as adversarial perturbations, into the original content. These signals aid in verifying the authenticity and integrity if the content is later altered. For effectiveness, these alterations visually resemble the original to evade detection while impairing the functionality of deepfake generation models. In the following subsections, we explore recent proactive detection efforts for GANs and DMs separately.
Figure 4: The comparison between passive and proactive deepfake detection methods. Figure from [154].
@@ 3.1.2.1   Proactive Methods for GANs
To further improve the resilience of perturbations, Huang et al. [153] introduce a poison perturbation generator that improves the effectiveness of perturbations using an alternating training strategy. Wang et al. [154] propose improving the resilience of perturbations by targeting the Lab color space, particularly the decorrelated a and b channels. Expanding defensive strategies, Anejaet al. [155] develop a data-driven method that quickly produces tailored perturbations to redirect manipulation attempts toward a predefined outcome. Huang et al. [156] design the Cross-Model Universal Adversarial Watermark, broadening defense across various images and deepfake models. Asnani et al. [157] and [158] propose using designed real image templates to improve the detection accuracy of manipulation, helping to differentiate between real and fake images.
Recent innovations have suggested conducting source tracing alongside injecting adversarial perturbations into the original input to enhance traceability against manipulations. Innovations by Wang et al. [159] and Wu et al. [160] underscore the importance of tracing original source images to boost detection system credibility. Yu et al. [14] enhance traceability by embedding traceable fingerprints directly into the generator’s parameters, linking outputs uniquely to their creators. Addressing gaps in proactive defenses, Tang et al. [161] introduce the Feature-Output ensemble UNiversal Disruptor, which disrupts in the latent space and employs a gradient-ensemble algorithm to provide a universal attack pattern against multiple deepfake models. In [162], Zhang et al. introduce Dual Defense, which embeds an invisible watermark in target faces to disrupt malicious face swapping while preserving traceability. This approach uses a watermark embedding network that balances invisibility, adversariality, and traceability through perceptual adversarial encoding strategies. In a recent study, Wu et al. [163] demonstrate that traditional watermarking methods, when applied to deepfake detection, can interfere with forgery detection. To address this issue, they propose AdvMark, a plug-and-play solution that retains the ability to extract watermarks for provenance tracking.
@@ 3.1.2.2   Proactive Methods for DMs
Expanding the scope, Zhao et al. [166] examine the effectiveness of watermarking within diffusion models, with Cao et al. [167] extending proactive security measures to audio models. To enhance model transferability, Shim et al. [168] disrupt the latent encoding phase of deepfake generation using Latent Ensemble ATtack, employing a Normalized Gradient Ensemble strategy to optimize various disruptions.
@@ 3.2 Multi-modal Deepfake Detection
The concurrence of fake content in multiple modalities leads to more convincing deepfakes compared to those visually manipulated ones. In such context, it becomes imperative to develop the capability to detect these intricate multi-modal deepfakes. In this subsection, we provide a comprehensive investigation of the more intricate and challenging multi-modal scenario, where auditory or textual content can also be falsified along with the presence of visual manipulation.
3.2.1 Audio-Visual Deepfake Detection
Audio-visual deepfakes involve the manipulation of either or both audio and visual streams in a video. The combination of manipulations on these two modalities potentially raises more serious ethical and social concerns, for example, manipulating politicians to make false claims convincingly by delivering a lip-synced video with fake audio.
The heterogeneous nature and the modality gap between audio and video make multi-modal audio-visual deepfake detection a more challenging task. Existing explorations on this task commonly design two-stream framework to incorporate multi-modal inputs. As shown in Figure 5, the efforts can be categorized into three main paradigms: independent learning, joint learning, and matching-based learning.
@@ 3.2.1 Audio-Visual Deepfake Detection
As a baseline method presented in [10], two independent streams of convolutional neural network are proposed separately for detecting audio and visual deepfakes, and the final result of multimodal detection is obtained by combining the audio and visual predictions. In a preliminary attempt to evaluate the FakeAVCeleb[5] benchmark, Khalid et al.[11] ensemble predictions from independently learned uni-modal detectors. Furthermore, Hashmi et al.[169] devise an audio-visual network besides uni-modal networks and adopt ensemble learning to make the final decision based on predictions from audio, video, and audio-visual networks. Similarly, AVFakeNet[170] leverages Swin-Transformer [197] as the backbones for both audio and visual branches for feature extraction and then integrates the result from both branches for the multi-modal prediction.
@@
Despite its intuitive nature, the paradigm of independent learning could overlook the intrinsic relationships between audio and visual data, which can be critical in accurately detecting deepfakes.
@@
Joint Learning This approach integrates audio and visual modalities to leverage multi-modal information for enhanced deepfake detection. The general network architecture of the joint learning paradigm encompasses two modal-specific branches taking as input corresponding modality content for feature extraction and a fusion module to integrate features of different modalities for audio-visual detection. The feature fusion module in this paradigm aids to mine the inter-modal relationship. Based on at which stage the modal-specific features are fused together, joint learning-based detection algorithms can be further divided into intermediate and late fusion manners. Moreover, some existing late fusion-based audio-visual detectors leverage multi-task strategy or impose regularization to facilitate the learned multi-modal representations.
@@
Intermediate Fusion: Under the assumption that simply combining the high-level abstract features extracted from both streams is insufficient to mine the inter-modal information, Zhou et al.[10] initially propose the joint audio-visual deepfake detection, handling cases in which either or both visual and auditory modalities have been manipulated. In their method, a sync-stream is designed to fuse the intermediate features from distinct modalities by cross-attention mechanism. Yang et al.[45] employ a Transformer architecture, where audio and visual inputs go through a temporal-spatial encoder, and corresponding features of both modalities interact with each other via a decoder with a bidirectional cross-attention module. Wang et al.[171] incorporate inspiration from how the biological system perceives the multi-modal environment. Marcella et al. [172] extract audio and visual features separately, followed by computing distance and attention maps between the audio and spatial visual features. To address the influence of irrelevant visual regions, such as hair and background, an attention map is applied to diminish their impact on the final classification. In this study, the intermediate audio and visual features attend to each other at all stages and are concatenated at the last few stages for further multi-modal detection.
@@
Late Fusion: This strategy only merges the audio and visual representations in the final stages of feature extraction. Techniques such as concatenation or addition are commonly used to merge modalities in late fusion [11][169][173][174][175][35][13]. Various advanced strategies, including the attention mechanism[176][177] and the MLP mixer layer[12] are applied to enhance effective fusion.
@@
Further, Liu et al. [175] propose a tri-branch framework to handle audio, video, and multi-scale frame inputs, where the extracted features are weighted based on inter-modal interactions and concatenated. Moreover, this method also designs Forgery Clues Magnification Transformer to improve feature extraction of each branch and Distribution Difference based Inconsistency Computing module to align information of audio and video modalities.
Dynamic weighted addition has proven effective in fusion strategies for multi-modal tasks. While addressing the temporal audio-visual deepfake localization task, Cai et al.[35] calculate the weighted sum of uni-modal boundary maps from corresponding branches as the fused boundary map, where the weights are dependent on uni-modal boundary maps themselves. Similarly, Wang et al.[13] leverage dynamic weighted averaging as late fusion approach in the proposed framework. They fuse modal-specific information via weighted averaging the classification tokens in both uni-modal features, which are extracted by a two-stream transformer. Notably, the weights in this method are calculated by inter-modal attention mechanism.
More sophisticated than simple concatenation or addition, attention mechanisms and MLP mixer layers offer advanced alternatives for multi-modal feature fusion. Sree et al.[177] employ a multi-head self-attention module after concatenating features of different modalities. The attention mechanism further enhances the inter-modal interaction and benefits subsequent multi-modal detection. More than the attention mechanism, Zou et al.[176] feed the concatenated audio-visual features into Transformer layers to fuse information from both modalities. Alternatively, Raza et al.[12] employ an MLP mixer layer that not only enhances intra-modal features but also facilitates the integration of inter-modal information by passing concatenated modal-specific features through an inter-modality MLP mixer layer.
@@
The networks proposed in [10, 45] are tasked with three binary classification objectives: uni-modal audio deepfake detection, uni-modal visual deepfake detection, and multi-modal detection based on the fused signals. Besides simultaneously introducing uni-modal detection tasks, Muppalla et al.[174] make the multi-modal detection more fine-grained, changing the original binary classification to a four-class one. Similarly, Raza et al.[12] reformulate multi-modal audio-visual deepfake detection as a multi-label classification task. They modify the classification head to incorporate fused signals and simultaneously perform classification tasks for both modalities. The multi-task strategy employed in these studies requires more fine-grained labels of each modality to provide supervision during training.
@@
Unlike those incorporating fine-grained uni-modal detection tasks, Yu et al.[178] design an auxiliary task from a disparate perspective, emphasizing learning content-based cross-modal correlations. In this work, Yu et al. introduce a correlation distillation task, where the student detection network learns to model cross-modal correlation from well-trained audio and visual speech recognition models. This explicitly learned cross-modal correlation based on content helps to prevent overfitting.
@@
Enhanced Representation: This line of work learns more discriminative representations for detection during training. To this end, various strategies including regularization[45][175][176] and disentanglement[177] are employed.
@@
Yang et al.[45] apply a contrastive loss and an additive angular margin loss to make the fused multi-modal features of manipulated and genuine samples more discriminative. Likewise, Zou et al.[176] also impose the contrastive constraint in cross-modal cases, thus the audio and visual features become more similar when they are both pristine. Besides, Zou et al.[176] further introduce within-modal regularization to align uni-modal features , ensuring the uni-modal features are discriminative for corresponding uni-modal detection. Liu et al.[175] also leverage contrastive loss [198] to impose regularization on the relationship between audio and video features based on Jensen-Shannon divergence [199].
@@
Mis-AvoiDD [177] improves multi-modal deepfake detection by decoupling representations into modal-invariant and modal-specific subspaces. It employs a modal-invariant encoder and two modal-specific encoders to extract common and specific features. This method minimizes discrepancies between modal-specific features in the invariant subspace and uses orthogonal loss to maintain distinctness between subspaces, ensuring effective feature alignment.
@@
Matching-based Learning Distinct from fusion-based learning, matching-based learning leverages the natural synchronization found in genuine audio-visual content, which is typically absent in synthetic pairs[179, 180]. This approach focuses on the degree to which the audio and visual streams consistently match, rather than using traditional classification techniques. In this paradigm, networks employ modality-specific feature extractors and are trained on authentic content via contrastive learning, enhancing their generalization capabilities across various manipulation.
@@
Cheng et al.[179] propose Voice-Face matching detection based on the widely verified observation of the homogeneity between one person’s face and voice. The detection network involves only voice and face feature extractors, which are trained by contrastive InfoNCE[200] loss. At inference, the predicted genuineness of the audio-visual pair is determined by the consistency of features from these two modalities. Similar to [179], Tian et al.[180] evaluate the cross-consistency between audio and visual features to compute the matching score, which is compared to a threshold to obtain multi-modal detection results. The method addresses the trade-off observed in deepfake generation between temporal identity consistency and the synchronization of face motion with audio-visual elements by incorporating temporal identity intra-consistency. This method merely required self-supervised learning on real videos by a contrastive learning. Furthermore, Marcella et al.[181] introduce a novel loss that estimates the distance between the 1st-order statistics of audio and visual feature distributions. They maximize this loss for fake data and minimize it for real data, complementing the feature-distance loss.
@@
Feng et al. [182] approach the task from an anomaly detection perspective, requiring training only on real videos in a self-supervised manner. Initially, an audio-visual network is trained to estimate inter-modal synchronization scores, used to extract heuristically designed features. An auto-regressive model then learns the distribution of these features. POI-Forensics [183] focuses on identity consistency rather than cross-modal inconsistency. In this work, Cozzolino et al. argue that manipulation perturbs the portrayed identity’s characteristics. This method requires a reference set of at least 10 videos of the person of interest. The similarity index between the target and reference videos is used to determine the prediction. Haq et al. [184] introduce biological signal emotion psychology to improve deepfake detection. Their neurosymbolic framework performs intra-modal emotion transition analysis and inter-modal reasoning based on extracted emotions from each modality. This framework does not require training on deepfake datasets, relying on psychology and emotion knowledge. Yu et al. [185] propose a two-stage framework for detecting audio-visual deepfakes. The Predictive Visual-audio Alignment Self-supervision stage uses augmented views of visual signals interacting with audio information to obtain VA representations. The designed Multi-modal Deepfake Detection stage concatenates these representations for multi-modal prediction, applying cross-entropy, contrastive, and auxiliary losses. AVFF [186] also employs a two-stage framework. In the first stage, audio-visual representation is learned in a self-supervised manner from real videos using the A2V and V2A networks. These modules are fine-tuned in the second stage, followed by a classifier network. The representation learned in the first stage is highly discriminative for deepfake detection, demonstrating the effectiveness of the proposed audio-visual representation learning paradigm.
3.2.2 Text-Visual Deepfake Detection
Recent advances in text-visual deepfake detection have introduced sophisticated methods to address manipulations on text and visual modalities. Figure 6 illustrates the framework for this multi-modal detection and grounding task, which is conducted in [6].
Figure 6: HAMMER[6] for text-visual deepfake detection and grounding. Figure from [6].